OpenStudy (mendicant_bias):

(Intro Analysis) I'm trying to find the sum of an infinite series by partial fraction decomposition. I'm able to perform the PFD correctly but then don't understand what I need to do to find the value of the series. Problem and working below.

OpenStudy (mendicant_bias):

\[\text{Use Partial Fraction Decomposition to find the sum of the series}\]\[\sum_{n=1}^{\infty}\frac{1}{n(n+3)}\]

OpenStudy (mendicant_bias):

\[\sum_{n=1}^{\infty}\frac{1}{n(n+3)}=\sum_{n=1}^{\infty}\frac{A}{n}+\frac{B}{(n+3)}=\sum_{n=1}^{\infty}A(n+3)+Bn\]

OpenStudy (mendicant_bias):

\[\text{Let n = -3.}\]\[A(-3+3)+B(-3)=1\]\[-3B = 1; \ \ \ B = -\frac{1}{3}.\] I don't remember the exact reasoning here, but there's some logic like since there are no n terms in the numerator, the term \[(A+B)n\]must be equal to zero. Thus, A = -B, and\[A = \frac{1}{3}\]

OpenStudy (mendicant_bias):

From here I can plug these coefficients back in and take their limits, which both go to zero-but how does that help me? This tells me nothing as far as I am aware.

OpenStudy (mendicant_bias):

@dan815

OpenStudy (mendicant_bias):

@amistre64

OpenStudy (amistre64):

your partial is flawed

OpenStudy (amistre64):

\[\frac{A}{n}+\frac{B}{m}\ne Am+Bn\]

OpenStudy (amistre64):

\[\frac{1}{nm}=\frac{A}{n}+\frac{B}{m}\implies 1=Am+Bn\]

OpenStudy (amistre64):

i see that you shortcuted that but made an accounting for it later on ...

OpenStudy (amistre64):

let n=0 to solver for A, since B(0)=0

OpenStudy (amistre64):

1 = (A+B)n +3A is one form of it if we distribute ... but that doesnt seem useful to me, unless n=0 :)

OpenStudy (mendicant_bias):

Yeah, in any case, I think I figured it out after you have the coefficients, just gotta make sure I can get a general term of a new sequence out of it

OpenStudy (mendicant_bias):

If you plug into the solved coefficients terms, n=1 and upwards:\[a_1 = 1/4\]\[a_2=1/10\]\[a_3=1/18\]\[a_4=1/28\]\[a_5=1/40\]

OpenStudy (mendicant_bias):

I see the pattern, just having trouble putting it into a general term, one min

OpenStudy (amistre64):

notice that when we reach n+3, we are subtracting terms that already exist i believe

OpenStudy (amistre64):

write it out as the A/n + B/(n+3) structure instead

OpenStudy (amistre64):

this is called a telescoping series

OpenStudy (mendicant_bias):

I'm not sure what you're talking about, without subtracting, the second fractions I get, in order of n up, are: -1/12, -1/15, /1-18 -1/21, etc

OpenStudy (amistre64):

\[B = -A\]right?

OpenStudy (mendicant_bias):

Yeah

OpenStudy (amistre64):

+A/1 - A/(1+3) +A/2 -A/(2+3) +A/3 -A/(3+3) +A/4 -A/(4+3) adding this up we get A/4 -A/(1+3) = 0 in fact every n greater than 3 has a negation to it .

OpenStudy (amistre64):

1 - 4 2 - 5 3 - 6 4 - 7 5 - 8 1 2 3 4 - 4 5 - 5 6 - 6 7 - 7 8 - 8 ...

OpenStudy (amistre64):

so the sum is just the first 3 terms, since they do not recieve a negating term

OpenStudy (amistre64):

spose set A=(1,2,3,4,5,6,...) set B=(4,5,6,...) A-B = (1,2,3) right?

OpenStudy (mendicant_bias):

Yeah, I see it, just had to write it out and perform the cancellations. The only remaining terms are A/1, A/2. and A/3.

OpenStudy (amistre64):

then thats the sum :)

OpenStudy (mendicant_bias):

How can I otherwise prove that or show that? While that makes sense it doesn't feel very rigorous

OpenStudy (amistre64):

proof it by a telescopic series.

OpenStudy (mendicant_bias):

Like the way with the 1/4, 1/10, 1/18, etc, it's 1/4, 1/4+6, 1/4+6+8, etc What?

OpenStudy (amistre64):

an inductive proof might be useful

OpenStudy (mendicant_bias):

Yeah, since this is analysis I feel like he won't let me go with, "they cancel out, so this is the sum", I'm not totally sure how he wants me to do this the way he wants

OpenStudy (amistre64):

show that for every term n>3, there is a cancelation. or are you wanting to develop a partial sum series?

OpenStudy (amistre64):

im not keen on methods tho

OpenStudy (dan815):

hi sry im here, r u done